Linear Systems
9. What volumes of 60% hydrochloric acid solution and 30% hydrochloric acid solution must be mixed to make 125 mL of 36% hydrochloric acid solution?




thank you very much!!!!

Answer by checkley77(12844)

.6x+.3(125-x)=.36*125
.6x+37.5-.3x=45
.3x=45-37.5
.3x=7.5
x=7.5/.3
x=25 ml of 60% acid is used.
125-25=100 ml of 30% acid is used.
Proof:
.6*25+.3*100=45
15+30=45
45=45